OpenStudy (lena772):

The functions f(x) and g(x) are shown below. f(x) = 0.05x g(x) = (0.05)^x Which statement best describes the graph of f(x) and g(x)? The graph of g(x) will eventually exceed the graph of f(x). The graph of f(x) will eventually exceed the graph of g(x). The graphs will both have their y-intercept equal to 1. The graphs will both have their y-intercept equal to 0.05

OpenStudy (lena772):

I believe we can eliminate C & D right off the bat because we don't know their y intercepts.

OpenStudy (kainui):

You can find their y-intercepts, since that's just when the graph is on the y-axis. When is it on the axis? When x=0. So plug in and find out! =D

OpenStudy (lena772):

f(x)=0

OpenStudy (lena772):

g(x)=1

OpenStudy (kainui):

Yeah, so now you can safely say C and D are garbage, just by looking at that f(x).

OpenStudy (lena772):

We can still eliminatE C and D

OpenStudy (lena772):

YUP

OpenStudy (usukidoll):

we could just plug in any x value and graph it... larger values produce wins

OpenStudy (lena772):

0.05^1=0.05 0.05*1=0.05

OpenStudy (kainui):

Yeah, I mean at this point look at the graph in your calculator until you can kind of visualize approximately how exponential and linear graphs behave differently.

OpenStudy (usukidoll):

try plug in 2 and beyond for x

OpenStudy (lena772):

0.05^2=0.0025 0.05*2=0.1

OpenStudy (lena772):

0.05^3=0.000125 0.05*3=0.15

OpenStudy (lena772):

must be B

OpenStudy (kc_kennylau):

bingo :D

OpenStudy (lena772):

But the graph you sent me makes it seem like the answer is A

OpenStudy (usukidoll):

but that graph isn't telling much

OpenStudy (kc_kennylau):

Well the graph also says B (sorry I gtg)

OpenStudy (usukidoll):

you have to really plug in a ton of values...like not just one because it doesn't say much..... Let x = 2,3...n for n is any number greater than 0

OpenStudy (usukidoll):

it's B

OpenStudy (usukidoll):

it's there on the graph, but very tiny